# Derived deformation functors, Koszul duality, and Maurer–Cartan spaces

### v1

@@ 3.1. Hinich’s simplicial nerve
with the obvious simplicial structure maps, where $\Omega^{\bullet}(\Delta^{n})$ is the cdga $\mathbb{Q}[t_{0},\ldots,t_{n},dt_{0},\ldots,dt_{n}]/(\sum t_{i}-1,\,\sum dt_{i})$ of de Rham polynomial forms on the $n$-simplex, with $t_{i}$ of degree 00.
@@
An argument similar to Lemma 2.4 ensures that $\underline{\mathfrak{Del}}(L,-)$ sends small extensions to fibrations of simplicial groupoids, and acyclic small extensions to trivial fibrations. Thus $\mathrm{diag}\,B\underline{\mathfrak{Del}}(L,-)$ is, like the other functors, fibrant in the model structure of Proposition 3.6 below, so we can check quasi-isomorphism on tangent spaces, where we have
$\pi_{i}\underline{\mathrm{MC}}(L,k[\epsilon_{n}])\cong\mathrm{H}^{n+1-i}(L)\cong\pi_{i}B\underline{\mathfrak{Del}}(L,k[\epsilon_{n}]).\qed$
@@
As in Example 1.10, the DGLA $L:=\underline{\mathrm{Der}}_{\mathcal{P}}(R,R)$ governs algebras $R$ over $k$-linear dg operads $\mathcal{P}$. We can now say that $\underline{\mathfrak{Del}}(L,A)$ is the simplicial groupoid whose objects are $A$-linear deformations of $R$, with degree $n$ isomorphisms between $\mathcal{P}\otimes A$-algebras $R^{\prime}$ and $R^{\prime\prime}$ given by $\mathcal{P}\otimes A\otimes\Omega^{\bullet}(\Delta^{n})$-algebra morphisms $R^{\prime}\otimes\Omega^{\bullet}(\Delta^{n})\to R^{\prime\prime}\otimes\Omega^{\bullet}(\Delta^{n})$ which are the identity modulo $\mathfrak{m}(A)$.
As in [Hin99], for $A\in dg_{+}\mathrm{Art}_{k}$ and $R$ cofibrant, with $\mathcal{P}$ and $R$ both concentrated in non-negative chain degrees, the simplicial set $\underline{\mathrm{MC}}(L,A)\simeq\mathrm{diag}\,B\underline{\mathfrak{Del}}(L,A)$ is weakly equivalent to the $\infty$-groupoid of derived deformations of $R$.
@@ 4.1. Representability and Schlessinger’s conditions
Given a functor $F\colon\thinspace dg_{+}\mathrm{Art}_{k}\to s\mathrm{Set}$, it is now natural to ask whether it is weakly equivalent to a fibrant object $G$ of $\mathrm{lex}(dg_{+}\mathrm{Art}_{k},s\mathrm{Set})$, i.e. whether there exists a zigzag of objectwise weak equivalences going from $F$ to $G$. It is fairly easy to see sufficient conditions on $F$, by identifying those properties of $G$ which are invariant under weak equivalence:
@@
Lurie [Lur11] refers to similar functors on weakly Artinian cdgas as formal moduli problems, but the conditions are neither sufficient nor necessary to endow a functor with a natural moduli interpretation, and Examples 1.9 and 1.10 mention examples of natural formal moduli problems which do not give rise to derived Schlessinger functors.
@@
in the homotopy category of simplicial sets.
@@
The following is then [Pri07, Theorem 4.14]. Essential surjectivity follows by applying Heller’s generalisation [Hel81] of Brown representability to $\pi_{0}\overrightarrow{F}$, along similar lines to Theorem 2.6.
@@ 4.2. The equivalences
Say that a map $f:A\to B$ in $DG^{+}dg_{+}\mathrm{Art}_{k}$ is a small extension if it is surjective with kernel $I$ satisfying $\mathfrak{m}(A)\cdot V=0$. Say that it is an acyclic small extension if moreover $\mathrm{H}_{*}(\mathrm{Tot}\,I)=0$.
@@ 5.1. Koszul duality
The proofs of the results of §2 rely only on Koszul duality between the Lie and non-unital commutative operads (via the non-unital algebras $\mathfrak{m}(A)$ associated to each local Artinian cdga $A\cong k\oplus\mathfrak{m}(A)$). They generalise to any dg operad $\mathcal{L}$ (generalising $\mathrm{Lie}$) and any dg co-operad $\mathcal{C}$ (generalising $\mathrm{Com}^{*}$) equipped with a morphism $\alpha\colon\thinspace\Omega\mathcal{C}\to\mathcal{L}$ in the notation of [LV12, §6.5.9] [^5] The datum $\alpha$ is known as an operadic twisting morphism, and itself arises as a Maurer–Cartan solution. for which the left and right twisted composite products $\mathcal{L}\circ_{\alpha}\mathcal{C}$ and $\mathcal{C}\circ_{\alpha}\mathcal{L}$ of [LV12, §6.4.11] are acyclic.
The map $\alpha$ ensures that for any $\mathcal{C}$-coalgebra $C$ and $\mathcal{L}$-algebra $L$, we have a DGLA structure on $\mathrm{Hom}_{k}(C,L)$, and hence a Maurer–Cartan set $\mathrm{MC}(\mathrm{Hom}(C,L))$. Under the additional conditions of [LV12, §6.6], the acyclicity conditions are equivalent to asking that $\alpha$ be a quasi-isomorphism.
@@ 5.1.1. The Quillen equivalence
Writing $\mathrm{DG}\mathcal{L}\mathrm{A}$ for category of $\mathcal{L}$-algebras in cochain complexes, $\mathrm{DG\mathcal{C}C}^{\mathrm{nilp}}$ for the category of ind-conilpotent $\mathcal{C}$-coalgebras in cochain complexes, i.e. coalgebras for the comonad $V\mapsto\bigoplus_{n}\mathcal{C}(n)\otimes V^{\otimes n}$, and $dg\mathrm{Art}(\mathcal{C})$ for the category of duals of finite-dimensional objects of $\mathrm{DG\mathcal{C}C}^{\mathrm{nilp}}$, we have a right Quillen equivalence
${\mathrm{DG}\mathcal{L}\mathrm{A}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{MC}}$$\scriptstyle{\beta}$${\mathrm{lex}(dg\mathrm{Art}(\mathcal{C}),\mathrm{Set})}$${\mathrm{DG\mathcal{C}C}^{\mathrm{nilp}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{pro}(dg\mathrm{Art}(\mathcal{C}))^{\mathrm{opp}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
for the projective model structure on $\mathrm{DG}\mathcal{L}\mathrm{A}$ (in particular with quasi-isomorphisms as weak equivalences) and a model structure “of the second kind” (by analogy with [Pos09]) on $\mathrm{DG\mathcal{C}C}^{\mathrm{nilp}}$ for which fibrant objects are those which are cofree as graded coalgebras (forgetting the differential), weak equivalences are $\beta^{*}$-quasi-isomorphisms, with generating cofibrations (resp. trivial cofibrations) dual to central (resp. acyclic central) extensions in $dg\mathrm{Art}(\mathcal{C})$. Weak equivalences between fibrant objects $C$ are maps inducing quasi-isomorphisms on complexes $\tan(C)\subset C$ of indecomposables666i.e. elements on which all operations vanish, a space isomorphic to the space of cogenerators.
@@ 5.1.2. Key features of the proof
The $\infty$-equivalence between $\mathcal{L}$-algebras up to quasi-isomorphism and $\mathcal{C}$-coalgebras up to $\beta^{*}$-quasi-isomorphism simply follows because the co-unit $\beta^{*}\beta L\to L$ of the adjunction is always a quasi-isomorphism, by acyclicity of $\mathcal{L}\circ_{\alpha}\mathcal{C}$ as in the proof of [LV12, Theorem 11.3.6]. A similar argument using the central series filtration shows that for fibrant $C\in\mathrm{DG\mathcal{C}C}^{\mathrm{nilp}}$, the tangent $\tan C\to\tan(\beta\beta^{*}C)\cong\beta^{*}C[1]$ of the unit is a quasi-isomorphism.
That $\beta^{*}$-quasi-isomorphisms are generated by duals $f\colon\thinspace C\to D$ of acyclic central extensions is more subtle. Centrality means the filtration $F_{0}D:=C$, $F_{1}D:=D$ on $D$ is compatible with the coalgebra structure, inducing an increasing filtration on $\beta^{*}D$ with associated graded $\beta^{*}(C\oplus\mathrm{coker\,}f)$, graded by powers of the acyclic complex $\mathrm{coker\,}f$, making $f$ a $\beta^{*}$-quasi-isomorphism. Conversely, as in the first part of proof of [LV12, Theorem 11.3.7], acyclicity of $\mathcal{C}\circ_{\alpha}\mathcal{L}$ implies the unit $C\to\beta\beta^{*}C$ of the adjunction is a transfinite composition of such maps.
@@ 5.1.3. Generalisations
The characteristic 00 hypothesis for the equivalence of §55.15.1.1 is only needed because we work with symmetric operads; a similar equivalence exists over any base field for non-symmetric operads such as the associative operad.
@@ 5.2. Relating Manetti’s functors with simplicial functors
The logical next step in the comparison, as covered in §44.2, is to discard the negative cochain degrees of $\mathcal{C}$-coalgebras (dually, the negative chain degrees of $\mathcal{C}$-algebras), recovering the same data by working with simplicial functors. For this approach to work in general, we have to assume that $\mathcal{C}$ is concentrated in non-negative cochain degrees. [^7] There is an obvious generalisation if we allow $\mathcal{C}$ to be the total complex of a co-operad in chain cochain complexes, but the simplicial functors become less manageable because the category of test objects changes with each simplicial degree.
@@ 5.2.1. Quillen equivalences
${\mathrm{lex}(dg\mathrm{Art}(\mathcal{C}),\mathrm{Set})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{Tot}\,_{*}}$${\mathrm{lex}(DG^{+}dg_{+}\mathrm{Art}(\mathcal{C}),\mathrm{Set})}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\mathrm{lex}(dg_{+}\mathrm{Art}(\mathcal{C}),s\mathrm{Set})}$$\scriptstyle{D_{*}}$${\mathrm{DG\mathcal{C}C}^{\mathrm{nilp}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\mathrm{dg_{+}DG^{+}\mathcal{C}C}^{\mathrm{nilp}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{Tot}\,}$$\scriptstyle{D}$${(\mathrm{DG^{+}\mathcal{C}C}^{\mathrm{nilp}})^{\Delta^{\mathrm{opp}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
where again these are all model structures of the second kind.
Trivial fibrations in $\mathrm{lex}(dg\mathrm{Art}(\mathcal{C}),\mathrm{Set})$ and $\mathrm{lex}(DG^{+}dg_{+}\mathrm{Art}(\mathcal{C}),\mathrm{Set})$ send central extensions to surjections, while fibrations send acyclic extensions (resp. $\mathrm{Tot}\,$-acyclic central extensions) to surjections. Trivial fibrations in $\mathrm{lex}(dg_{+}\mathrm{Art}(\mathcal{C}),s\mathrm{Set})$ map central extensions to trivial Kan fibrations, while fibrations map central extensions to Kan fibrations and acyclic central extensions to trivial Kan fibrations.
Fibrant objects in $C\in\mathrm{dg_{+}DG^{+}\mathcal{C}C}^{\mathrm{nilp}}$ are those which are cofree as bigraded coalgebras and satisfy the additional condition that the subcomplex $\tan(C)\subset C$ of indecomposables satisfies $\mathrm{H}_{i}\tan(C)^{n}=0$ for all $i\geq 0,n\geq 0$ and $\mathrm{H}^{n}\tan(C)_{i}=0$ for all $i>0,n>0$, with reasoning similar to [Pri07, Lemma 1.56]. Morphisms between fibrant objects are weak equivalences precisely when they induce isomorphisms on $\mathrm{H}_{*}\mathrm{H}^{*}\tan$.

### v2

@@ 3.1. Hinich’s simplicial nerve
with the obvious simplicial structure maps, where $\Omega^{\bullet}(\Delta^{n})$ is the cdga $\mathbb{Q}[t_{0},\ldots,t_{n},dt_{0},\ldots,dt_{n}]/(\sum t_{i}-1,\,\sum dt_{i})$ of de Rham polynomial forms on the $n$-simplex, with $t_{i}$ of degree 0.
@@
The functor $\underline{\mathfrak{Del}}(L,-)$ preserves limits, and an argument similar to Lemma 2.4 ensures that it sends small extensions to fibrations of simplicial groupoids, and acyclic small extensions to trivial fibrations. For any simplicial set $X$ and any of the functors $F$ in Definition 3.2, the functor $X,F(-)]:=\mathrm{Hom}_{\mathrm{Ho}(s\mathrm{Set})}(X,F(-))$ is thus a deformation functor, so weak equivalence follows by checking isomorphism on tangent cohomology for spheres, where we have
$S^{i},\underline{\mathrm{MC}}(L,k[\epsilon_{n}])]\cong\mathrm{H}^{n}\mathrm{Def}(L\otimes\Omega^{\bullet}(S^{i}))\cong\mathrm{H}^{n+1}(L)\times\mathrm{H}^{n+1-i}(L)\cong[S^{i},B\underline{\mathfrak{Del}}(L,k[\epsilon_{n}])].\qed$
@@
As in Example 1.10, the DGLA $L:=\underline{\mathrm{Der}}_{\mathcal{P}}(R,R)$ governs deformations of algebras $R$ over $k$-linear dg operads $\mathcal{P}$. We can now say that $\underline{\mathfrak{Del}}(L,A)$ is the simplicial groupoid whose objects are $A$-linear deformations of $R$, with degree $n$ isomorphisms between $\mathcal{P}\otimes A$-algebras $R^{\prime}$ and $R^{\prime\prime}$ given by $\mathcal{P}\otimes A\otimes\Omega^{\bullet}(\Delta^{n})$-algebra morphisms $R^{\prime}\otimes\Omega^{\bullet}(\Delta^{n})\to R^{\prime\prime}\otimes\Omega^{\bullet}(\Delta^{n})$ which are the identity modulo $\mathfrak{m}(A)$.
As in [Hin99], for $A\in dg_{+}\mathrm{Art}_{k}$ and $R$ cofibrant, with $\mathcal{P}$ and $R$ both concentrated in non-negative chain degrees, the simplicial set $\underline{\mathrm{MC}}(L,A)\simeq\mathrm{diag}\,B\underline{\mathfrak{Del}}(L,A)$ is weakly equivalent to the $\infty$-groupoid of derived deformations of $R$. Without those constraints, it gives contraderived deformations as in [Pri25].
@@ 4.1. Representability and Schlessinger’s conditions
Given a functor $F\colon\thinspace dg_{+}\mathrm{Art}_{k}\to s\mathrm{Set}$, it is now natural to ask whether it is weakly equivalent to a fibrant object $G$ of $\mathrm{lex}(dg_{+}\mathrm{Art}_{k},s\mathrm{Set})$, i.e. whether there exists a zigzag of objectwise weak equivalences going from $F$ to $G$. It is fairly easy to see necessary conditions on $F$, by identifying those properties of $G$ which are invariant under weak equivalence:
@@
Lurie [Lur11] refers to similar functors on weakly Artinian cdgas as formal moduli problems, but the conditions are neither sufficient nor necessary to endow a functor with a natural moduli interpretation, and Examples 1.9 and 1.10 mention examples of natural formal moduli problems (in the traditionally understood sense) which do not give rise to derived Schlessinger functors (i.e. formal moduli problems in Lurie’s sense).
@@
in the homotopy category of simplicial sets, where $I[1]$ is square-zero.
@@
Remark 4.4.
As for instance in [Pri10, Definition 1.13], we can define tangent cohomology groups [^7] The notation $\mathrm{D}$ is based on that for André–Quillen cohomology groups in [Qui68], with which these coincide for representable functors. for derived Schlessinger functors by $\mathrm{D}^{n-i}(F,V):=\pi_{i}F(k\oplus V[n])$ for $i,n\geq 0$ and $V$ a finite-dimensional non-negatively graded $k$-chain complex, equipped with zero multiplication.
The derived Schlessinger conditions ensure that this is well-defined and also imply that $\mathrm{D}^{n}(F,V)\cong\bigoplus_{j}\mathrm{D}^{n+j}(F,k)\otimes\mathrm{H}_{j}V$. Repeated application of Proposition 4.3 implies that tangent cohomology groups detect weak equivalences.
The following is then [Pri07, Theorem 4.14]; although only stated there on the level of homotopy categories, it implies full faithfulness on simplicial localisation because the functor preserves the cotensoring of 3.8, so for $K\in s\mathrm{Set}$ we have
$\mathrm{Hom}_{\mathrm{Ho}(s\mathrm{Set})}(K,\mathbf{R}\mathrm{map}_{\mathrm{lex}(dg_{+}\mathrm{Art}_{k},s\mathrm{Set})_{\mathrm{fib}}}(F,G))\cong\pi_{0}\mathbf{R}\mathrm{map}_{\mathrm{lex}(dg_{+}\mathrm{Art}_{k},s\mathrm{Set})_{\mathrm{fib}}}(F,G^{K})$
$\cong\pi_{0}\mathbf{R}\mathrm{map}_{\mathcal{S}}(F,G^{K})\cong\mathrm{Hom}_{\mathrm{Ho}(s\mathrm{Set})}(K,\mathbf{R}\mathrm{map}_{\mathcal{S}}(F,G)).$
Essential surjectivity is established there by applying Heller’s generalisation [Hel81] of Brown representability to $\pi_{0}\overrightarrow{F}$, along similar lines to Theorem 2.6.
@@
See [Lur11, Theorem 0.0.13] for a later variant in the setting of Remark 1.3, previously stated without proof as [Lur04, Remark 6.2.5].
@@ 4.2. The equivalences
Say that a map $f:A\to B$ in $DG^{+}dg_{+}\mathrm{Art}_{k}$ is a small extension if it is surjective with kernel $I$ satisfying $\mathfrak{m}(A)\cdot I=0$. Say that it is an acyclic small extension if moreover $\mathrm{H}_{*}(\mathrm{Tot}\,I)=0$.
@@ 5.1. Koszul duality
The proofs of the results of §2 rely only on Koszul duality between the Lie and non-unital commutative operads (via the non-unital algebras $\mathfrak{m}(A)$ associated to each local Artinian cdga $A\cong k\oplus\mathfrak{m}(A)$). They generalise to any dg operad $\mathcal{L}$ (generalising $\mathrm{Lie}$) and any dg co-operad $\mathcal{C}$ (generalising $\mathrm{Com}^{*}$) equipped with a morphism $\alpha\colon\thinspace\Omega\mathcal{C}\to s\mathcal{L}$ in the notation of [LV12, §6.5.9] [^8] The datum $\alpha$ is known as an operadic twisting morphism, and itself arises as a Maurer–Cartan solution. for which the twisted composite products $\mathcal{L}\circ_{\alpha}\mathcal{C}$ and $\mathcal{C}\circ_{\alpha}\mathcal{L}$ of [LV12, §6.4.11] are acyclic.
The map $\alpha$ ensures that for any $\mathcal{C}$-coalgebra $C$ and $\mathcal{L}$-algebra $L$, we have a set $\mathrm{MC}(\mathrm{Hom}(C,L))$ as in [LV12, Definition 11.1.1]. Under the additional conditions of [LV12, §6.6], the acyclicity conditions are equivalent to asking that $\alpha$ be a quasi-isomorphism.
@@ 5.1.1. The Quillen equivalence
Writing $\mathrm{DG}\mathcal{L}\mathrm{A}$ for category of $\mathcal{L}$-algebras in cochain complexes, $\mathrm{DG\mathcal{C}C}^{\mathrm{nilp}}$ for the category of ind-conilpotent $\mathcal{C}$-coalgebras in cochain complexes, i.e. coalgebras for the comonad $V\mapsto\bigoplus_{n}\mathcal{C}(n)\otimes_{k[S_{n}]}V^{\otimes n}$, and $dg\mathrm{Art}(\mathcal{C})$ for the category of duals of finite-dimensional objects of $\mathrm{DG\mathcal{C}C}^{\mathrm{nilp}}$, by [Val14, Theorem 2.1] we have a right Quillen equivalence $\beta$:
${\mathrm{DG}\mathcal{L}\mathrm{A}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{MC}}$$\scriptstyle{\beta}$${\mathrm{lex}(dg\mathrm{Art}(\mathcal{C}),\mathrm{Set})}$${\mathrm{DG\mathcal{C}C}^{\mathrm{nilp}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${\mathrm{pro}(dg\mathrm{Art}(\mathcal{C}))^{\mathrm{opp}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces,}\textstyle{\sim}$
for the projective model structure on $\mathrm{DG}\mathcal{L}\mathrm{A}$ (in particular with quasi-isomorphisms as weak equivalences) and a model structure “of the second kind” (by analogy with [Pos09]) on $\mathrm{DG\mathcal{C}C}^{\mathrm{nilp}}$ for which fibrant objects are those which are cofree as graded coalgebras (forgetting the differential) and weak equivalences are $\beta^{*}$-quasi-isomorphisms. Weak equivalences between fibrant objects $C$ are maps inducing quasi-isomorphisms on complexes $\tan(C)\subset C$ of indecomposables, i.e. elements on which all operations vanish (isomorphic to cogenerators).
We can go further and say that generating cofibrations (resp. trivial cofibrations) in $\mathrm{DG\mathcal{C}C}^{\mathrm{nilp}}$ are dual to central (resp. acyclic central) extensions in $dg\mathrm{Art}(\mathcal{C})$.
@@ 5.1.2. Key features of the proof
The $\infty$-equivalence in [Val14] between $\mathcal{L}$-algebras up to quasi-isomorphism and $\mathcal{C}$-coalgebras up to $\beta^{*}$-quasi-isomorphism simply follows because the co-unit $\beta^{*}\beta L\to L$ of the adjunction is always a quasi-isomorphism, by acyclicity of $\mathcal{L}\circ_{\alpha}\mathcal{C}$. A similar argument using the central series filtration shows that for fibrant $C\in\mathrm{DG\mathcal{C}C}^{\mathrm{nilp}}$, the tangent $\tan C\to\tan(\beta\beta^{*}C)\cong\beta^{*}C[1]$ of the unit is a quasi-isomorphism.
That $\beta^{*}$-quasi-isomorphisms are generated by duals $f\colon\thinspace C\to D$ of acyclic central extensions. is more subtle. Centrality means the filtration $F_{0}D:=C$, $F_{1}D:=D$ on $D$ is compatible with the coalgebra structure, inducing an increasing filtration on $\beta^{*}D$ with associated graded $\beta^{*}(C\oplus\mathrm{coker\,}f)$, graded by powers of the acyclic complex $\mathrm{coker\,}f$, making $f$ a $\beta^{*}$-quasi-isomorphism. Conversely, as in the first part of proof of [LV12, Theorem 11.3.7], acyclicity of $\mathcal{C}\circ_{\alpha}\mathcal{L}$ implies the unit $C\to\beta\beta^{*}C$ of the adjunction is a transfinite composition of such maps.
@@ 5.1.3. Generalisations
The characteristic 0 hypothesis for the equivalence of §55.15.1.1 is only needed because we work with symmetric operads; a similar equivalence exists over any base field for non-symmetric operads such as the associative operad.
@@ 5.2. Relating Manetti’s functors with simplicial functors
The logical next step in the comparison, as covered in §44.2, is to discard the negative cochain degrees of $\mathcal{C}$-coalgebras (dually, the negative chain degrees of $\mathcal{C}$-algebras), recovering the same data by working with simplicial functors. For this approach to work in general, we have to assume that $\mathcal{C}$ is concentrated in non-negative cochain degrees. [^9] There is an obvious generalisation if we allow $\mathcal{C}$ to be the total complex of a co-operad in chain cochain complexes, but the simplicial functors become less manageable because the category of test objects changes with each simplicial degree.  [^10] However, comparison with weakly Artinian objects as in Remark 1.3 relies on coconnectivity of the Lie operad to ensure an analogue of [Boo20, Lemma 4.3.1], specifically that for all $i$, $\mathrm{H}^{i-n}\mathcal{L}(n)=0$ for $n\gg 0$.
@@ 5.2.1. Quillen equivalences
${\mathrm{lex}(dg\mathrm{Art}(\mathcal{C}),\mathrm{Set})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{Tot}\,_{*}}$${\mathrm{lex}(DG^{+}dg_{+}\mathrm{Art}(\mathcal{C}),\mathrm{Set})}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\mathrm{lex}(dg_{+}\mathrm{Art}(\mathcal{C}),s\mathrm{Set})}$$\scriptstyle{D_{*}}$${\mathrm{DG\mathcal{C}C}^{\mathrm{nilp}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\sim}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\mathrm{dg_{+}DG^{+}\mathcal{C}C}^{\mathrm{nilp}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{Tot}\,}$$\scriptstyle{D}\textstyle{\sim}$${(\mathrm{DG^{+}\mathcal{C}C}^{\mathrm{nilp}})^{\Delta^{\mathrm{opp}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces,}\textstyle{\sim}$
where weak equivalence in these model structures is again stronger than quasi-isomorphism.
Trivial fibrations in $\mathrm{lex}(dg\mathrm{Art}(\mathcal{C}),\mathrm{Set})$ and $\mathrm{lex}(DG^{+}dg_{+}\mathrm{Art}(\mathcal{C}),\mathrm{Set})$ send central extensions to surjections, while fibrations send acyclic extensions (resp. $\mathrm{Tot}\,$-acyclic central extensions) to surjections. Trivial fibrations in $\mathrm{lex}(dg_{+}\mathrm{Art}(\mathcal{C}),s\mathrm{Set})$ map central extensions to trivial Kan fibrations, while fibrations map central extensions to Kan fibrations and acyclic central extensions to trivial Kan fibrations.  [^11] One step in establishing the model structure [Pri07, Theorem 2.14] on $\mathrm{lex}(dg_{+}\mathrm{Art}(\mathcal{C}),s\mathrm{Set})$ requires a more refined argument outside the commutative setting. Given a central extension $A\to B$ with kernel $M$, and a cofibration $A\to S$, the first unnumbered corrigendum lemma preceding that theorem uses the description of the pushout as a central extension with kernel $M\otimes_{A}S$. In general, we instead have to consider the filtration $F$ on $S$ generated by $M\subset F^{1}S$, so $\mathrm{gr}_{F}^{0}S\cong S\coprod_{A}B$ and $\hat{\mathrm{gr}}_{F}S\cong(\mathrm{gr}_{F}^{0}S)\coprod_{B}(B\oplus M)$. Then $S$ is the limit of a sequence of abelian extensions $S/F^{p+1}S\to S/F^{p}S$, with associated obstructions in $\mathrm{H}^{2}\mathrm{Tot}\,^{\Pi}\underline{\mathrm{Der}}(\mathrm{gr}_{F}^{0}S,\mathrm{gr}_{F}^{p}S)$. When $M$ has a contracting homotopy, it is inherited by $\mathrm{gr}_{F}^{p}S$ for $p>0$, so the obstruction vanishes.
Fibrant objects $C\in\mathrm{dg_{+}DG^{+}\mathcal{C}C}^{\mathrm{nilp}}$ are those which are cofree as bigraded ind-conilpotent coalgebras and satisfy the additional condition that the subcomplex $\tan(C)\subset C$ of indecomposables satisfies $\mathrm{H}_{i}\tan(C)^{n}=0$ for all $i\geq 0,n\geq 0$ and $\mathrm{H}^{n}\tan(C)_{i}=0$ for all $i>0,n>0$, with reasoning similar to [Pri07, Lemma 1.56]. Morphisms between fibrant objects are weak equivalences precisely when they induce isomorphisms on $\mathrm{H}_{*}\mathrm{H}^{*}\tan$.